# Unsupervised Learning for Quadratic Assignment

### v1

@@ 5.2 Cross-size Generalization
Our model naturally generalizes across problem sizes due to its permutation-equivariant design, where the soft permutation matrix $\mathbb{T}\in\mathbb{R}^{n\times n}$ automatically adapts to match input dimensions. Experiments show that a model trained on $n=100$ effectively generalizes to $n=200$ problems and vice versa, while consistently outperforming random initialization, as shown in Table 8 and Table 9. For instance, when a model trained on $n=200$ is applied to $n=100$ problems, UL initialization achieves a solution cost of 1795.57 compared to 1805.83 for random initialization, and when used with TS$(1k,25,25)$, PLUME TS reaches 1688.43 versus 1692.78 for standard TS with random initialization.
@@ 6 Conclusion
In this paper, we propose PLUME search, a framework to enhance combinatorial optimization through UL. By leveraging permutation-based loss, we demonstrate that neural networks can effectively learn QAP directly from instances instead of using SL/RL. Our experimental results suggest that UL can generate high-quality initial solutions that significantly outperform random initialization and these better starting points consistently lead to better final solutions after tabu search. This is achieved by a simple NN structure with negligible computational overhead. Our approach demonstrates strong generalization across varying problem densities and sizes.
PLUME search takes a different approach from traditional heuristics and other SL/RL methods, offering a complementary method that integrates seamlessly with existing frameworks rather than competing with them. Future work might explore extending this approach to other permutation-based combinatorial problems, developing more sophisticated neural architectures to capture problem structure, and investigating the theoretical foundations of UL for combinatorial optimization.

### v2

@@ 5.2 Cross-size Generalization
Our model naturally generalizes across problem sizes due to its permutation-equivariant design, where the soft permutation matrix $\mathbb{T}\in\mathbb{R}^{n\times n}$ automatically adapts to match input dimensions. Experiments show that a model trained on $n=100$ effectively generalizes to $n=200$ problems and vice versa, while consistently outperforming random initialization, as shown in Table 5. For instance, when a model trained on $n=200$ is applied to $n=100$ problems, UL initialization achieves a solution cost of 1754.69 compared to 1805.83 for random initialization, and when used with TS(1,000,25,25)(1{,}000,25,25), PLUME TS reaches 1674.92 versus 1692.78 for standard TS with random initialization.
@@ 6 Conclusion
In this paper, we propose PLUME search, a framework to enhance combinatorial optimization through unsupervised learning. By leveraging a permutation-based loss, we demonstrate that neural networks can effectively learn the quadratic assignment problem directly from instances, rather than relying on supervised or reinforcement learning. Our experimental results indicate that UL can generate high-quality initial solutions that significantly outperform random initialization, and these improved starting points consistently lead to superior final solutions after tabu search, while also exhibiting strong generalization across varying problem densities and sizes.
A key insight is that our method does not merely act as an initializer, but rather learns to capture and exploit the intrinsic problem structure. In this sense, our unsupervised method can itself be regarded as a form of search, providing solutions of comparable quality to conventional heuristics after substantial exploration. PLUME search therefore takes a different path from traditional heuristics, offering a complementary paradigm that integrates seamlessly with existing frameworks rather than competing with them.